INRARED SPECTROSCOPY AND TUNNELING DYNAMICS OF THE VINYL RADICAL IN ^{4}He NAN-ODROPLETS

PAUL L. RASTON, TAO LIANG, \underline{EMMANUEL I. OBI}, and GARY E. DOUBERLY, Department of Chemistry, University of Georgia, Athens, Georgia 30602, USA.

The vinyl radical has been trapped in ^{4}He nanodroplets and probed with infrared laser spectroscopy in the CH stretch region between 2850 and 3200 cm^{−1}. The assigned band origins for the CH_{2} symmetric (ν_{3}), CH_{2} antisymmetric (ν_{2}), and lone α-CH stretch (ν_{1}) vibrations are in good agreement with previously reported full-dimensional vibrational configuration interaction calculations.^{a} For all three bands, a-type and b-type transitions are observed from the lowest symmetry allowed roconvibrational state of each nuclear spin isomer, which allows for a determination of the tunneling splittings in both the ground and excited vibrational levels. Comparisons to gas phase millimeter-wave rotation-tunneling^{b} and high-resolution jet-cooled infrared spectra^{c} reveal that the effect of the ^{4}He solvent is to reduce the ground and ν_{3} excited state tunneling splittings by ≈20%. This solvent-induced modification of the tunneling dynamics can be reasonably accounted for by assuming either an ≈2.5% increase in the effective barrier height along the tunneling coordinate or an ≈5% increase in the effective reduced mass of the tunneling particles.

aA. R. Sharma, B. J. Braams, S. Carter, B. C. Shepler, and J. M. Bowman, J. Chem. Phys. 130, 174301 2009

bK. Tanaka, M. Toshimitsu, K. Harada, and T. Tanaka, J. Chem. Phys. 120, 3604 2004

cF. Dong, M. Roberts, and D. J. Nesbitt, J. Chem. Phys. 128, 044305 (2008)